OpenStudy (anonymous):

If a+b+c= 0, then find the value of (a+b-c)^3 + (b+c-a)^3 + (c+a-b)^3

OpenStudy (kirbykirby):

I don't know if this helps much but: If a+b+c=0, then a+b=-c and a+c=-b and b+c=-a So, (a+b-c)^3 = (-c-c)^3=-8c^3 by substitution "a+b+ with the above (b+c-a)^3=(-a-a)^3=-8a^3 (c+a-b)^3=(-b-b)^3=-8b^3 => the expression becomes -8(a^3+b^3+c^3)

OpenStudy (kirbykirby):

I'm not sure what else to do

OpenStudy (kirbykirby):

What class is this for?

OpenStudy (kirbykirby):

Is there more information

OpenStudy (anonymous):

If a, b, and c equal 0, then no matter what you do to it, it will always be 0 right?

OpenStudy (kirbykirby):

^It says a+b+c=0, not a=b=c=0, so you could have like a=-1,b=1,c=0 or a=-2,b=-1,c=3, etc.

OpenStudy (shubhamsrg):

a^3 + b^3 + c^3 - 3abc = (a+b+c)(a^2+ b^2 + c^2 - ab -bc -ca) You may use this :

OpenStudy (aravindg):

did u get the answer?

OpenStudy (anonymous):

no..

OpenStudy (aravindg):

why not ..where are u stuck?

OpenStudy (anonymous):

i m not getting ..i m stuck how 2 start

OpenStudy (aravindg):

use the property @shubhamsrg told ..its really easy after that

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

i m not getting it..

OpenStudy (anonymous):

a^3 + b^3 + c^3 - 3abc = (a+b+c)(a^2+ b^2 + c^2 - ab -bc -ca). How to use this.

OpenStudy (shubhamsrg):

a+b+c =0 make that substitution on RHS

OpenStudy (anonymous):

a^3 + b^3 + c^3 - 3abc = (a+b+c)(a^2+ b^2 + c^2 - ab -bc -ca). a^3 + b^3 + c^3 - 3abc = (0)(a^2+ b^2 + c^2 - ab -bc -ca) a^3 + b^3 + c^3 - 3abc = 0

OpenStudy (anonymous):

a^3 + b^3 + c^3 = 3abc

OpenStudy (anonymous):

vat after that

OpenStudy (shubhamsrg):

substitute that into @kirbykirby 's work.

OpenStudy (anonymous):

=> the expression becomes= -8(a^3+b^3+c^3) => the expression becomes= -8(3abc)=-24abc

OpenStudy (anonymous):

is that answer.

OpenStudy (shubhamsrg):

Yep, seems right.

OpenStudy (anonymous):

thanks